VIBRATIONAL PREDISSOCIATION SPECTRA IN THE SHARED PROTON REGION OF PROTONATED FORMIC ACID WIRES

\underline{HELEN GERARDI}, CHRIS LEAVITT, and MARK JOHNSON, Yale University, Department of Chemistry, New Haven, CT.

Protonated formic acid networks are thought to form long chains where the excess charge is distributed along the embedded H-bonds holding it together. This system therefore effectively captures anhydrous proton transport at an intermediate state, thereby allowing a detailed look at the underlying mechanism. We present vibrational predissociation spectra of the Ar-tagged (HCOOH)_{n}H^{+} cluster ions, formed via supersonic expansion of argon gas seeded with formic acid. The resulting data reveal the first well-defined spectral signatures associated with the collective vibrational motions of the H-bonding backbone, and we follow how these features evolve as the number of formic acid molecules is increased from 2 to 5. Bands in the 700-1500 cm^{−1} range are found to strongly depend on cluster size, and the shifts displaying an even-odd alternation before stabilizing in the n=4 and 5 clusters. We interpret these trends in the context of charge delocalization and the resulting effect on the carbonyl stretches.